OpenStudy (anonymous):

convert the rectanguler equation to a polar equation. Express r in terms of theta. y-x=8 so is it just x^2-y^2=64 r=8? but thats not theta....

OpenStudy (anonymous):

you ve done it wrong y = r cos(theta) x = r sin(theta)

OpenStudy (anonymous):

y - x = 8 r cos(theta) - r sin(theta) = 8 r (cos(theta) - sin(theta)) = 8 r = 8/(cos(theta)-sin(theta))

OpenStudy (anonymous):

thank you